OpenStudy (anonymous):

How do you simplify y^2-9/3+9y? Thanks :)

OpenStudy (anonymous):

What? How do you get -y?

OpenStudy (anonymous):

Yes. I want to understand it. I get a different answer...

OpenStudy (anonymous):

Is it supposed to be 3y+9?

OpenStudy (anonymous):

The original problem

OpenStudy (anonymous):

\[y^2-9=(y+3)(y-3)\]\[3y+9=3(y+3)\] The y+3 will cancel out, so you get \[(y-3)/3\]

OpenStudy (anonymous):

but the problem says 3+9y.. :/

OpenStudy (anonymous):

Hmm...that's not going to simplify nicely.

OpenStudy (anonymous):

-y won't be correct

OpenStudy (anonymous):

Yeah its going to be hard. I dont know what to do..

OpenStudy (anonymous):

It can't be simplified in any meaningful way. Are you sure the problem is written correctly?

OpenStudy (anonymous):

I got it in my college algebra book so it should be correct? Unless its one of those trick questions.

OpenStudy (anonymous):

Probably...jerks. Is it supposed to be simplified by synthetic division?

OpenStudy (anonymous):

No. It just says simplify the following rational expressions. :/

OpenStudy (anonymous):

Ah, in that case, it's simplified already.

OpenStudy (anonymous):

But cant you simplify the denominator into 3(1+3y)??

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

You can do that, but it's not really any simpler.

OpenStudy (anonymous):

what are you gonna do with that ?

OpenStudy (anonymous):

Oh ok. Well thank you!! :D

OpenStudy (anonymous):

Thats what I was thinking.. Is it?

OpenStudy (anonymous):

a-b looks good to me.

OpenStudy (anonymous):

You have to assume that a is larger than or equal to b